OpenStudy (anonymous):

I Will Medal And Fan! I need explanation AND answer! A group of 8 students was asked, "How many hours did you watch television last week?" Here are their responses. 7, 10, 12, 17, 9, 16, 14, 12 Find the mean number of hours for these students. If necessary, round your answer to the nearest tenth.

pooja195 (pooja195):

Add up all those numbers first

pooja195 (pooja195):

7+10=?

OpenStudy (anonymous):

17

pooja195 (pooja195):

17+12=?

OpenStudy (anonymous):

@pooja195 knows you add the numbers and divide by the number of people that watched TV

pooja195 (pooja195):

@BPDlkeme234 i know.

OpenStudy (anonymous):

29

pooja195 (pooja195):

29+17

OpenStudy (anonymous):

46

pooja195 (pooja195):

46+9=?

OpenStudy (anonymous):

@TheForgottenDogs sum up all the numbers given

OpenStudy (anonymous):

7+10+12+17+9+16+14+12

OpenStudy (anonymous):

Then divide by the total amount of numbers you have 1=7 2=10 3=12 etc I count 8 numbers

OpenStudy (anonymous):

so 7+10 + 12 +17+ 9 + 6 +14 + 12 = 87

pooja195 (pooja195):

Why are you doing the work for them?

OpenStudy (anonymous):

Example, giving an example

OpenStudy (anonymous):

I'm sorry, I only now realised you are a qualified helper, I will exit here!

pooja195 (pooja195):

lol that means nothing u did a good job but it would have been better if u let the asker do the work next time so that they understand :)

OpenStudy (anonymous):

my sincere apologies @pooja195

pooja195 (pooja195):

its ok :) you did a good job!

OpenStudy (anonymous):

I don't get it, i'll just look it up

OpenStudy (anonymous):

@TheForgottenDogs what is the average of 2 apples and 5 apples?

OpenStudy (anonymous):

@TheForgottenDogs it is the toal number of apples divided by the number of apples

OpenStudy (anonymous):

The formula for the average = total number of objects (or numbers)/total amount of objects(numbers)